OpenStudy (anonymous):

The numbers 1-8 are written on a sheet of paper and the eight sheets of paper are placed in a bowl. If one sheet of paper is selected at random from the bowl, determine the probability that the number selected is odd

OpenStudy (anonymous):

P(E)=4/8=1/2